OpenStudy (hatahoooochee):

Help, I'll fan and medal Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = Four divided by x squared. + 9

OpenStudy (hatahoooochee):

Answers: f(x) = x + 9, g(x) = Four divided by x squared. f(x) = x, g(x) = Four divided by x. + 9 f(x) = One divided by x., g(x) = Four divided by x. + 9 f(x) = Four divided by x squared., g(x) = 9

OpenStudy (aveline):

Could you use the equation tool to write the question and answers? That way confusion can be prevented

OpenStudy (hatahoooochee):

Can you help? @Vocaloid I only have one more question after this one and I'll give you medals for both questions

OpenStudy (hatahoooochee):

\[A. f(x)=x+9, g(x)=4/x^2\]

OpenStudy (hatahoooochee):

\[B. f(x)=x g(x)=4/x +9\]

OpenStudy (hatahoooochee):

\[C. f(x)=1/x g(x)= 4/x+9\]

OpenStudy (hatahoooochee):

\[D. f(x)=4/x^{2} g(x)=9\]

Vocaloid (vocaloid):

basically, you want to write the function where g(x) can be inserted as x into f(x)

OpenStudy (aveline):

And \[y = \frac{ 4 }{ x^ {2} }+9\]?

OpenStudy (hatahoooochee):

I did that, but I'm still super confused

OpenStudy (aveline):

For example, in letter D, you would substitute 9^2 for x^2. We can cancel out D since it would equal 4/81. Does this make sense?

OpenStudy (hatahoooochee):

I think so

OpenStudy (hatahoooochee):

Yes, so then answer choice D is wrong?

OpenStudy (aveline):

D is incorrect.

OpenStudy (aveline):

Now let's look at A. Can you figure out whether it is right or wrong?

OpenStudy (hatahoooochee):

I think it's wrong

OpenStudy (aveline):

Substitute 4/x^2 into the f(x) equation. What do you get?

OpenStudy (hatahoooochee):

Ugh..I'm not sure, how do I figure it out again?

OpenStudy (aveline):

You plug in the value of the g(x) equation into the equation of f(x). For example, f(x)=1 + x and g(x)=4 If we were to simplify this, we would get 5 since we would substitute the g(x) value (a.k.a four) into the equation on the left.

OpenStudy (hatahoooochee):

So for A it would be 9/4?

OpenStudy (aveline):

\[A.f(x)=x+9, g(x)=4/x^ {2} \] Let's plug in g(x) into the first equation \[f(x)=g(x)+9\] \[f(x)=\frac{ 4 }{ x^ {2} }+9\]

OpenStudy (aveline):

Does that make sense?

OpenStudy (hatahoooochee):

I think so

OpenStudy (hatahoooochee):

Yes, it does

OpenStudy (hatahoooochee):

@Aveline ??

OpenStudy (aveline):

...we've got the answer, don't we?

OpenStudy (hatahoooochee):

Yes, if I open a new question, can you help with my last one and I'll give you another medal? You're a life saver.

OpenStudy (hatahoooochee):

I just need you to check my answer

OpenStudy (aveline):

I'm sorry, but I'm in computer class right now and it ends in 3 minutes >-< I'd love to help but I'm sure someone else will be willing to help you too :)

OpenStudy (hatahoooochee):

Okay, thanks anyways